OpenStudy (anonymous):

There are two consecutive even numbers such that six times the smaller number minus five times the greater number is -2. What is the smaller number? Answers: 10 12 8 14

OpenStudy (anonymous):

assume that the smaller number is x then the larger number l b x+2

OpenStudy (anonymous):

Let x=smaller number. From the question we know the following: \[6x-(5(x+2))=-2\]From here you just need to solve for x and make sure it's even. I'll give you a hint to start, start by distributing the 5 through the parentheses.

OpenStudy (anonymous):

\[(6*2n)+5(2n+2)=-2\]

OpenStudy (anonymous):

nw ur question says that six times the smaller number minus five times the greater number is -2 which means \[6x-5(x+2)=-2\] solve for x and x l b the ans

OpenStudy (anonymous):

\[(6*2n)-5(2n+2)=-2\]